OpenStudy (anonymous):

The percent composition of methyl butanoate is 58.8% C, 9.8% H, and 31.4 % O and its molar mass is 102 g/mol. What is its empirical formula? What is its molecular formula? @mayankdevnani

OpenStudy (mayankdevnani):

Assume you have 100g and you can convert % to g. m(C) = 58.8g m(H) = 9.8g m(O) = 31.4g n(C) = m/M = 58.8g / 12g/mol = 4.9mol n(H) = m/M = 9.8g / 1g/mol = 9.8mol n(O) = m/M = 31.4g / 16g/mol = 2.0mol These ratios are C:H:O 5:10:2 (after rounding off) So the empirical formula is C5H10O2 This compound has a molecular mass of: m=nM = 12x5 + 1x10 + 16x2 = 60+10+32 = 102g/mol

OpenStudy (mayankdevnani):

got it @malutella

OpenStudy (mayankdevnani):

ok @malutella

OpenStudy (anonymous):

Thank youu!

OpenStudy (mayankdevnani):

welcome